Vegetable prices in Pakistan are skyrocketing. Tomato prices in local markets
have shot up to Rs. 220-250 per kg while prices of other vegetables have also
increased considerably. According to official rates, the price of tomatoes was
Rs180 per kg but its price has touched Rs. 300 in retail markets which led to a
lower quantity demanded of this vegetable. The given table describes what
happened to prices and the quantities of tomatoes demanded.
Price
(Rupees)
Quantity Demanded
(Kilograms)
200 800
300 750
Requirements:
A. Calculate price elasticity of demand for tomatoes when price is between
Rs. 200 and Rs. 300 by using arc elasticity method.
B. After calculation, interpret your result whether demand for tomatoes is
elastic or inelastic.
1
Expert's answer
2019-11-28T11:38:10-0500

Solution:

A)

Let "p_1=200, p_2=300, Q_1=800, Q_2=750" .


"E= \frac {Q_2-Q_1}{P_2-P_1} \times \frac {P_2+P_1}{Q_2+Q_1}"



"E= \frac {750-800}{300-200} \times \frac {300+200}{750+800} = - \frac {11}{31}"

B)

"\vert E \vert = \vert - \frac {11}{31} \vert < 1"

Therefore, the demand for tomatoes is not elastic.